Solve the following system of linear equations using (a) matrix inversion, and (b) Cramer's rule. [ 1 3 ox] To 0.5 1 y 10.5 0 1Z
1
Expert's answer
2020-05-19T09:02:07-0400

We are given that "A =\begin{bmatrix}
 1 & 3 & 0 \\
 0 & 0.5 & 1 \\
 0.5 & 0 & 1
\end{bmatrix}", "b = \begin{bmatrix}
 4 \\
 1 \\
 4
\end{bmatrix}"

(a) Using the matrix inversion

"A^{-1} =\frac{1}{det(A)}A^{-1}=\frac{1}{1*0.5*1 + 3*0.5*0.5} \begin{bmatrix}
 0.5 & 0.5 & -0.25 \\
 -3 & 1 & 1.5 \\
 3 & -1 & 0.5
\end{bmatrix}^{T} = \begin{bmatrix}
 0.25 & -1.5 & 1.5 \\
 0.25 & 0.5 & -0.5 \\
 -0.125 & 0.75 & 0.25
\end{bmatrix}"

"\begin{bmatrix}
 x \\
 y \\
 z
\end{bmatrix} = A^{-1} b = \begin{bmatrix}
 5.5 \\
 -0.5 \\
 1.25
\end{bmatrix}"


(b) Using the Cramer's rule

"det(A_1)=det\begin{bmatrix}
 4 & 3 & 0 \\
 1 & 0.5 & 1 \\
 4 & 0 & 1
\end{bmatrix} = 4 * 0.5 * 1 + 3 * 1 * (-1 + 4) = 11"

"det(A_2)=det\begin{bmatrix}
 1 & 4 & 0 \\
 0 & 1 & 1 \\
 0.5 & 4 & 1
\end{bmatrix} = 1 * (1 * 1 - 1 * 4) + 4 * 1 * 0.5 = -1"


"det(A_3)=det\begin{bmatrix}
 1 & 3 & 4 \\
 0 & 0.5 & 1 \\
 0.5 & 0 & 4
\end{bmatrix} = 1 * 0.5 * 4 + 3 * 1 * 0.5 - 4 * 0.5 * 0.5 = 2.5"

"\begin{bmatrix}
 x \\
 y \\
 z
\end{bmatrix} = \begin{bmatrix}
 det(A_1)/det(A) \\
 det(A_2)/det(A) \\
 det(A_3)/det(A)
\end{bmatrix} = \begin{bmatrix}
 5.5 \\
 -0.5 \\
 1.25
\end{bmatrix}"